OpenStudy (anonymous):

A parallelogram has sides measuring 23.8 m and 35.3 m. The height corresponding to the 23.8-m base is 9.9 m. Find the height, to the nearest tenth of a meter, corresponding to the 35.3-m base. A. 8317.4 m B. 14.7 m C. 6.7 m D. none of these

OpenStudy (nurali):

the area of parallelogram is : side times height corresponding to that side, so 23,8*9,9=35,3*h h=6.6764 Correct answer is C. 6.7

OpenStudy (anonymous):

Thank you !

OpenStudy (nurali):

My Pleasure.